OpenStudy (anonymous):

f(x) = x^2 – 4 and g(x) = x – 2 (a) What is (f + g)(x)? (b) What is (f – g)(x)? (c) What is (fg)(x)? (d) What is (f/g)(x)

OpenStudy (dinamix):

f+g = x^2+x-6 f-g= x^2 -x-2 fg = (x^2-4)(x-2) = x^3-4x-2x^2+8 = x^3-2x^2-4x+8 f/g = x^2-4/x-2 = (x-2)(x+2)/(x-2) =x+2

rebeccaxhawaii (rebeccaxhawaii):

got it?

jhonyy9 (jhonyy9):

@dinamix not directly answers please !!!! try teaching the asker not give the right answers never please !!!!